OpenStudy (anonymous):

How would I go about solving this? I'm lost? Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10

OpenStudy (anonymous):

linear algebra?

OpenStudy (anonymous):

System of equations with three variables.

OpenStudy (saifoo.khan):

z = -10 -x -4y <---- eq1.

OpenStudy (saifoo.khan):

3x -3y +6( -10 -x -4y) = -10

OpenStudy (saifoo.khan):

-3 x-27 y-50 = 0

OpenStudy (saifoo.khan):

same as with second one, eliminate "z" from the 3rd equation as well.

OpenStudy (anonymous):

let me try and illl post my trials...

OpenStudy (anonymous):

If I remember correctly you can apply the Gram Schmit process and come up with the same

OpenStudy (saifoo.khan):

x + 2y + 2z = –10 x + 2y + 2(-10 -x -4y) = –10 -x-6 y = 10

OpenStudy (saifoo.khan):

now, we have: -3x-27y-50 = 0 -x - 6y = 10

OpenStudy (saifoo.khan):

Solve for x and y.

OpenStudy (anonymous):

(1, –2, –3) (–4, –1, –2) (2, –2, –4) (–3, –2, –3) those are my answers

OpenStudy (anonymous):

possible answers btw

OpenStudy (saifoo.khan):

B

OpenStudy (anonymous):

okay, but how do i get b?

OpenStudy (saifoo.khan):

Calc. ;)

OpenStudy (saifoo.khan):

im trying to figure it out..

OpenStudy (anonymous):

but i have to show my work :O

OpenStudy (anonymous):

thanks :)

OpenStudy (saifoo.khan):

im doing it in my book, wait please.

OpenStudy (saifoo.khan):

Got it. xD

OpenStudy (saifoo.khan):

x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10 Solving eq 1 for z: z = -10 - x -4y Now, subsitute this value in both eq 2 and eq 3 3x – 3y + 6z = –21 & x + 2y + 2z = –10 3x -3y + 6(-10 - x -4y) = -21 x + 2y +2(-10 - x -4y) = -10 Solve and you will get; Solve and you will get; -3x - 27y = 39 -x -6y = 10

OpenStudy (saifoo.khan):

now solve these two for x and y. so x = -4, y = -1 as now you have both, solve them in any equation for z. And you are done!! :)

OpenStudy (phi):

Here's how people do it using linear algebra. It tends to get to the right answer without to much confusion.